# Avoiding Catastrophic Forgetting in Visual Classification Using Human Concept Formation

### v1

@@ Abstract
Deep neural networks have excelled in machine learning, particularly in vision tasks, however, they often suffer from catastrophic forgetting when learning new tasks sequentially. In this work, we propose Cobweb/4V, a novel visual classification approach that builds on Cobweb, a human-like learning system that is inspired by the way humans incrementally learn new concepts over time. In this research, we conduct a comprehensive evaluation, showcasing Cobweb/4V’s proficiency in learning visual concepts, requiring less data to achieve effective learning outcomes compared to traditional methods, maintaining stable performance over time, and achieving commendable asymptotic behavior, without catastrophic forgetting effects. These characteristics align with learning strategies in human cognition, positioning Cobweb/4V as a promising alternative to neural network approaches. Keywords: human-like learning, concept formation, continual learning, information theory
@@ Introduction
Computer vision has made substantial progress in recent years due to breakthroughs in deep neural networks. Neural networks can exceed human capabilities in certain tasks such as object detection and classification (?, ?). However, such networks cannot handle continual learning of new tasks without forgetting previously learned data. Catastrophic forgetting is a fundamental challenge for artificial neural networks (?, ?). This phenomenon happens when the network is trained on multiple tasks sequentially, and to meet the objective of the new task, it changes the weights learned to perform the previous tasks. To prevent it, the model needs to be trained on both new data and previous data simultaneously, which consumes more time and energy, underscoring the trade-off between efficiency and performance in neural networks.
@@
As humans, we excel at learning knowledge and skills across different domains, adapting them through new experiences, and transferring them throughout our lifespan (?, ?, ?). While we gradually forget some of our previously learned knowledge over time, it rarely happens that our existing knowledge catastrophically fades away when learning something new. To achieve more efficiency, while maintaining high performance, neural networks need new human-like capabilities for incrementally learning new concepts over time (?, ?). Over the past two decades, extensive research has been dedicated to studying catastrophic forgetting, aiming to mitigate its impact on neural networks. These approaches can be categorized into four main groups, 1) regularization-based, 2) replay-based, 3) optimization-based, and 4) architecture-based methods (?, ?). While contributing significantly to mitigating this phenomenon, these methods come with limitations including reliance on memory-intensive mechanisms, difficulties in preserving old parameters when dealing with a growing number of tasks, and sensitivity to task-specific parameters.
@@
In this work we present Cobweb/4V, a novel visual classification approach that builds upon Cobweb (?, ?, ?), a concept formation technique that draws inspiration from the psychological study of how humans learn concepts. Cobweb supports all of the human-like learning capabilities outlined in (?, ?), such as learning rapidly in an incremental, piecemeal way, while explaining basic-level and typicality effects from psychology (?, ?). This work aims to combine computer vision principles with this prior approach and explore the idea of incorporating new visual information incrementally without erasing previously learned data. Our results demonstrate that Cobweb/4V does not exhibit catastrophic forgetting, only limited interference effects when compared to neural networks. We find that Cobweb/4V is competitive with neural network approaches while having minimal forgetting effects. It is also more data efficient and achieves asymptotic performance with fewer examples. Overall, Cobweb/4V’s capacity to preserve previously acquired knowledge while incrementally incorporating new experiences makes it a promising alternative to neural networks. Moreover, it provides insights into learning mechanisms that closely resemble human processing, particularly when applied to vision tasks.
Related Works
Continual, Lifelong Learning
@@
Continual learning refers to the ability of a learning model to adapt to new tasks without forgetting previously learned knowledge. It is an important area of research in machine learning as it helps machine learning models to become more human-like by being able to accumulate knowledge over their lifetime, improves data efficiency, and enhances generalization capabilities by applying past knowledge to new, unseen tasks, and more importantly, addresses the challenge of catastrophic forgetting, where models tend to lose previously learned information when exposed to new tasks. Catastrophic forgetting happens due to a trade-off between learning plasticity and memory stability. Earlier efforts to explore continual learning (?, ?) laid the foundation for addressing catastrophic interference in connectionist networks, highlighting the challenges of sequential learning and the need for adaptive algorithms in machine learning. Since then, extensive research has been dedicated to addressing this fundamental challenge, which can be categorized into four main categories, the replay-, optimization-, regularization-, and architecture-based methods (?, ?).
@@
Replay-based methods tackle catastrophic forgetting by revisiting data from past experiences during the learning of new tasks. Early approaches, such as rehearsal in backpropagation networks (?, ?), underscored the advantage of revisiting previously learned patterns. Following that, the development of the pseudo-rehearsal technique (?, ?) was prompted by scenarios in which old training data was not accessible, which led to the adoption of a memory network for generating pseudo-examples during the retraining. The Generative Adversarial Networks (GANs) (?, ?) led to a new approach called deep generative replay (?, ?), which employs a dual-model architecture with a generator based on GANs for generating “fake” inputs from the previous training data distribution and a solver for image classification. However, replay methods often involve revisiting past examples, which increases computational costs and training time. In addition, capturing the diversity of past experiences can be challenging, with generative replay struggling to reproduce a wide range of learning scenarios, potentially leading to incomplete learning and reduced performance.
@@
Optimization-based methods offer a new paradigm to address the limitations associated with replay, by adjusting the optimization program. Gradient Episodic Memory (GEM) (?, ?) is an example that integrates samples from earlier tasks into the loss function, using an episodic memory to store these samples, so it effectively balances adapting to new information and preserving knowledge from prior tasks without requiring access to previous models. However, GEM relies on task boundaries during training and experiences performance degradation with reduced memory size. As more memories accumulate, the optimization problem becomes more constrained, negatively impacting knowledge transfer to later tasks.
@@
Regularization-based methods enhance a model’s adaptation to new tasks while maintaining previous knowledge by adding regularization penalties and constraints to the training loss function to penalize significant changes to crucial parameters. The weight regularization method preserves critical parameters for prior tasks to reduce catastrophic forgetting, such as Elastic Weight Consolidation (EWC) (?, ?), which estimates the importance of each weight regarding previous tasks using the Fisher information matrix. The function regularization method targets a function’s intermediate or final output, such as Learning without Forgetting (LwF) (?, ?), which reduces forgetting by introducing a regularization term and computes the distillation loss (?, ?).
@@
Architecture-based methods involve designing model architectures that inherently support continual learning. One example is the Progressive Neural Networks (PNNs) (?, ?) which employs modular networks, and adds new columns or neural networks for each subsequent task while maintaining existing ones. This ensures the forward transfer of learned features from one network to another, preventing interference with learning without forgetting. However, as tasks increase, a PNN’s complexity grows, leading to higher computational demands and inefficiencies in memory and computational resources (?, ?). Additionally, PNNs lack positive backward transfer, meaning knowledge from newer tasks cannot enhance performance on earlier tasks (?, ?).
@@
While continual learning is a pivotal step towards achieving human-like learning, it encounters various challenges, including the widespread issue of catastrophic forgetting that extends across different approaches such as regularization and optimization-based methods (?, ?, ?). Additionally, issues related to scalability and computational cost are evident in replay- and architecture-based methods (?, ?, ?). In our evaluation of Cobweb/4V, we draw insights from a recent study (?, ?) that extensively explored continual learning approaches and bring attention to their limitations, particularly in scenarios involving the learning of new classes, notably with regularization-based methods. To align our evaluation with best practices, we choose to benchmark Cobweb/4V against a replay-based approach, highlighting the widely recognized effectiveness of replay in shaping successful continual learning strategies.
Cobweb
@@
This approach presents an alternative to neural networks, supporting incremental and unsupervised learning over a stream of examples (?, ?, ?). Given consecutive instances, Cobweb learns a hierarchical concept tree in a fashion inspired by how humans form concepts. Each instance is represented by a set of discrete attribute-value pairs (e.g. {color: blue; shape: circle; size: large}), and each concept node in a Cobweb tree is represented by a table of attribute-value count pairs of its adopted instances (e.g. {color: {red: 1, blue: 2}, shape: {circle: 1, square: 2}, size: {small: 2, large: 1}}).
Given an instance (not necessarily represented with all specified attribute values), Cobweb sorts it downward through its hierarchy: Starting at the root node, in every branch (a concept node with its children), Cobweb recursively considers whether to direct the instance to one of the children or whether to stop at the current node. When this process terminates, it can use the final concept’s representation to predict the value(s) of its unobserved attribute(s).
@@
When learning, Cobweb follows a similar approach to prediction, but as it categorizes the instance down its tree it updates the counts of concepts along the categorization path to reflect the attribute values of the new instance. At each branch point, Cobweb also considers three additional restructuring operations: merging (combining the two “best” children and reconsidering the available operations), splitting (removing the “best” child and elevating its children to the current level), and creating (creating a new child node which adopts the new learned instance only). During both prediction and learning, Cobweb uses category utility (?, ?) to guide its choices, which evaluates the increase in the predictive power of a child node compared to its parent. In each categorization and learning step, Cobweb proceeds with the operation that maximizes category utility. While the original Cobweb approach only supports nominal attribute values, Cobweb/3 (?, ?) extends this support to continuous attributes by using normal distributions to describe their probabilities. As a result, each concept stores the mean and standard deviation for each continuous attribute rather than the attribute value counts, which are maintained for nominal values.
Cobweb/4V
@@
This paper introduces Cobweb/4V, a new Cobweb variant designed for visual learning tasks. It extends a prior implementation by ? (?) with a new prediction approach that uses multiple concepts throughout the categorization tree to yield improved predictive performance. It also introduces a new tensor-based concept representation that enables efficient processing of images.
Learning with Mutual Information
@@
All prior Cobweb studies employed the probabilistic category utility (?, ?) in both predicting and learning processes, which calculates the expected increase in the ability to predict attribute values correctly given the knowledge of an instance’s membership in a particular concept. In Cobweb/4V, however, we use the information-theoretic category utility (?, ?). This involves establishing a connection between feature predictability and informativeness. Using this new measure, the category utility of a concept $c$ is given by the following equation:
$CU(c)=P(c)[U(c_{p})-U(c)]$ (1)
Here $c_{p}$ is the parent of $c$ and $U(c)$ is the uncertainty (or entropy) of $c$. This measure, which is equivalent to mutual information, indicates the amount of information gained by knowledge of the child class over the parent. It is the unsupervised equivalent of the information gain measure used with decision trees (?, ?).
Predicting with a Combination of Concepts
@@
During prediction, Cobweb first finds a final concept at the subordinate level in a greedy way, then uses the count information associated with this node to predict the unobserved attribute values of the categorized instance. This approach has been widely adopted in prior studies because it seems to yield good empirical performance (?, ?, ?, ?). However, other studies have explored alternative concept levels for predictions, such as using basic-level concepts (?, ?, ?).
@@
Our revised approach introduces a new prediction strategy that combines predictions from multiple concepts within the tree. Given an instance $\bm{x}$ with some unobserved feature(s) and a maximum number of the concepts to use for prediction ($N_{max}$), Cobweb begins categorization at the root. However, instead of sorting $\bm{x}$ down a single, greedy path, Cobweb/4V instead uses best-first search. At each point in the search, it expands the node $c^{*}$ on the search frontier that has the greatest score $s(c)=P(c|\bm{x})P(\bm{x}|c)$. This heuristic, which is sometimes referred to as collocation (?, ?), [^1] Prior work shows it is related to probabilistic category utility; the expected correct guesses of a concept can be reinterpreted as a form of weighted collocation (?, ?). is the product of cue and category validity. Once a $c^{*}$ is identified, the system adds it to an expanded node list. Next, the system evaluates the collocation of the children of $c^{*}$ and adds these to the search frontier to be considered for expansion. This process repeats until Cobweb/4V has expanded $N_{max}$ nodes. If $\mathcal{C}^{*}$ denotes all the nodes expanded during categorization, Cobweb/4V predicts the attribute $X_{i}$ will have value $x_{i}$ with probability:
@@
This represents the combination of predictions from all expanded nodes, weighted by their collocation (the softmax ensures the sum of the weights to one). Although Cobweb/4V only expands $N_{max}$ nodes, rather than all the nodes, this approach is effectively a form of Bayesian model averaging (?, ?).
A New Tensor Representation
In Cobweb/4V, each instance is represented by a tensor of pixels and a class label, as opposed to a list of attribute-value pairs, as used in previous Cobweb approaches. This adaptation better enables it to support vision learning tasks, where each instance is an image represented as a tensor of pixel intensities. Similar to Cobweb/3, each concept in Cobweb/4V stores the means and standard deviations of values for each tensor index; however, to enable more efficient processing, it also stores these statistics in a tensor representation.
Similar to Cobweb/3, the values for each attribute are assumed to be conditionally independent given the concept and to follow a normal distribution. Consequently, the uncertainty of a node $c$ is the sum of entropy values for the attributes under $c$, each treated as an independent normal distribution.
In a vision classification task, instances are represented by tensors, where each attribute corresponds to pixel intensity from the image. Additionally, each instance is associated with its ground-truth classification value, leading each concept to maintain counts for each potential classification value based on the adopted instances—using the standard attribute-value representation for labels as Cobweb normally does. The use of tensor representations for instances (our implementation uses PyTorch) significantly enhances speed, yielding substantially faster processing than versions of Cobweb that utilize attribute-value lists.
Experiments
We evaluate Cobweb/4V and two neural network baselines across distinct experimental conditions with the MNIST dataset (60,000 images in the training set, 10,000 images in the test set), prevalently used in the continual learning studies (?, ?, ?). [^2] The codes for the experiments are available at
https://github.com/Teachable-AI-Lab/cobweb-vision
Preliminary Experiment

### v2

# Incremental Concept Formation over Visual Images Without Catastrophic Forgetting

@@ Abstract
Deep neural networks have excelled in machine learning, particularly in vision tasks, however, they often suffer from catastrophic forgetting when learning new tasks sequentially. In this work, we introduce Cobweb/4V, an alternative to traditional neural network approaches. Cobweb/4V is a novel visual classification method that builds on Cobweb, a human-like learning system that is inspired by the way humans incrementally learn new concepts over time. In this research, we conduct a comprehensive evaluation, showcasing Cobweb/4V’s proficiency in learning visual concepts, requiring less data to achieve effective learning outcomes compared to traditional methods, maintaining stable performance over time, and achieving commendable asymptotic behavior, without catastrophic forgetting effects. These characteristics align with learning strategies in human cognition, positioning Cobweb/4V as a promising alternative to neural network approaches.
nicki.barari@drexel.edu
Drexel University, Philadelphia, PA 19104 USA xinlian@gatech.edu cmaclell@gatech.edu
Teachable AI Lab, Georgia Institute of Technology, Atlanta, GA 30332 USA
@@ 1 Introduction
Computer vision has made substantial progress in recent years due to breakthroughs in deep neural networks. Neural networks can exceed human capabilities in certain tasks such as object detection and classification (He et al., 2016). However, such networks cannot handle continual learning of new tasks without forgetting previously learned data. Catastrophic forgetting is a fundamental challenge for artificial neural networks (McCloskey & Cohen, 1989). This phenomenon happens when the network is trained on multiple tasks sequentially, and to meet the objective of the new task, it changes the weights learned to perform the previous tasks. To prevent it, the model needs to be trained on both new data and previous data simultaneously, which consumes more time and energy, underscoring the trade-off between efficiency and performance in neural networks.
@@
As humans, we excel at learning knowledge and skills across different domains, adapting them through new experiences, and transferring them throughout our lifespan (Barnett & Ceci, 2002; Calvert et al., 2004). While we gradually forget some of our previously learned knowledge over time, it rarely happens that our existing knowledge catastrophically fades away when learning something new. To achieve more efficiency, while maintaining high performance, neural networks need new human-like capabilities for incrementally learning new concepts over time (French, 1999). Over the past two decades, extensive research has been dedicated to studying catastrophic forgetting, aiming to mitigate its impact on neural networks. These approaches can be categorized into four main groups, 1) regularization-based, 2) replay-based, 3) optimization-based, and 4) architecture-based methods (Wang et al., 2023). While contributing significantly to mitigating this phenomenon, these methods come with limitations including reliance on memory-intensive mechanisms, difficulties in preserving old parameters when dealing with a growing number of tasks, and sensitivity to task-specific parameters.
@@
In this work we present Cobweb/4V, a novel visual classification approach that builds upon Cobweb (Fisher et al., 2014; Gennari et al., 1989), a concept formation technique that draws inspiration from the psychological study of how humans learn concepts. Cobweb supports all of the human-like learning capabilities outlined in (Langley, 2022), such as learning rapidly in an incremental, piecemeal way, while explaining basic-level and typicality effects from psychology (Fisher, 1988). This work aims to combine computer vision principles with this prior approach and explore the idea of incorporating new visual information incrementally without erasing previously learned data. Our results demonstrate that Cobweb/4V does not exhibit catastrophic forgetting, only limited interference effects when compared to neural networks. We find that Cobweb/4V is competitive with neural network approaches while having minimal forgetting effects. It is also more data efficient and achieves asymptotic performance with fewer examples. Overall, Cobweb/4V’s capacity to preserve previously acquired knowledge while incrementally incorporating new experiences makes it a promising alternative to neural networks. Moreover, it provides insights into learning mechanisms that closely resemble human processing, particularly when applied to vision tasks.
2 Related Works
2.1 Continual, Lifelong Learning
@@
Continual learning refers to the ability of a learning model to adapt to new tasks without forgetting previously learned knowledge. It is an important area of research in machine learning as it helps machine learning models to become more human-like by being able to accumulate knowledge over their lifetime, improves data efficiency, and enhances generalization capabilities by applying past knowledge to new, unseen tasks, and more importantly, addresses the challenge of catastrophic forgetting, where models tend to lose previously learned information when exposed to new tasks. Catastrophic forgetting happens due to a trade-off between learning plasticity and memory stability. Earlier efforts to explore continual learning (McCloskey & Cohen, 1989) laid the foundation for addressing catastrophic interference in connectionist networks, highlighting the challenges of sequential learning and the need for adaptive algorithms in machine learning. Since then, extensive research has been dedicated to addressing this fundamental challenge, which can be categorized into four main categories, the replay-, optimization-, regularization-, and architecture-based methods (Wang et al., 2023).
@@
Replay-based methods tackle catastrophic forgetting by revisiting data from past experiences during the learning of new tasks. Early approaches, such as rehearsal in backpropagation networks (Robins, 1993), underscored the advantage of revisiting previously learned patterns. Following that, the development of the pseudo-rehearsal technique (Robins, 1995) was prompted by scenarios in which old training data was not accessible, which led to the adoption of a memory network for generating pseudo-examples during the retraining. The Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) led to a new approach called deep generative replay (Shin et al., 2017), which employs a dual-model architecture with a generator based on GANs for generating “fake” inputs from the previous training data distribution and a solver for image classification. However, replay methods often involve revisiting past examples, which increases computational costs and training time. In addition, capturing the diversity of past experiences can be challenging, with generative replay struggling to reproduce a wide range of learning scenarios, potentially leading to incomplete learning and reduced performance.
@@
Optimization-based methods offer a new paradigm to address the limitations associated with replay, by adjusting the optimization program. Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017) is an example that integrates samples from earlier tasks into the loss function, using an episodic memory to store these samples, so it effectively balances adapting to new information and preserving knowledge from prior tasks without requiring access to previous models. However, GEM relies on task boundaries during training and experiences performance degradation with reduced memory size. As more memories accumulate, the optimization problem becomes more constrained, negatively impacting knowledge transfer to later tasks.
@@
Regularization-based methods enhance a model’s adaptation to new tasks while maintaining previous knowledge by adding regularization penalties and constraints to the training loss function to penalize significant changes to crucial parameters. The weight regularization method preserves critical parameters for prior tasks to reduce catastrophic forgetting, such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), which estimates the importance of each weight regarding previous tasks using the Fisher information matrix. The function regularization method targets a function’s intermediate or final output, such as Learning without Forgetting (LwF) (Li & Hoiem, 2017), which reduces forgetting by introducing a regularization term and computes the distillation loss (Hinton et al., 2015).
@@
Architecture-based methods involve designing model architectures that inherently support continual learning. One example is the Progressive Neural Networks (PNNs) (Rusu et al., 2016) which employs modular networks, and adds new columns or neural networks for each subsequent task while maintaining existing ones. This ensures the forward transfer of learned features from one network to another, preventing interference with learning without forgetting. However, as tasks increase, a PNN’s complexity grows, leading to higher computational demands and inefficiencies in memory and computational resources (Parisi et al., 2019). Additionally, PNNs lack positive backward transfer, meaning knowledge from newer tasks cannot enhance performance on earlier tasks (Schwarz et al., 2018).
@@
While continual learning is a pivotal step towards achieving human-like learning, it encounters various challenges, including the widespread issue of catastrophic forgetting that extends across different approaches such as regularization and optimization-based methods (Kirkpatrick et al., 2017; Li & Hoiem, 2017). Additionally, issues related to scalability and computational cost are evident in replay- and architecture-based methods (Robins, 1995; Rusu et al., 2016). In our evaluation of Cobweb/4V, we draw insights from a recent study (Van de Ven & Tolias, 2019) that extensively explored continual learning approaches and bring attention to their limitations, particularly in scenarios involving the learning of new classes, notably with regularization-based methods. To align our evaluation with best practices, we choose to benchmark Cobweb/4V against a replay-based approach, highlighting the widely recognized effectiveness of replay in shaping successful continual learning strategies.
2.2 Cobweb
@@
This approach presents an alternative to neural networks, supporting incremental and unsupervised learning over a stream of examples (Fisher, 1987; Fisher et al., 2014). Given consecutive instances, Cobweb learns a concept hierarchy in a fashion inspired by how humans form concepts. Each instance is represented by a set of discrete attribute-value pairs, e.g. {color: blue; shape: square}, and each concept node in a Cobweb tree is represented by a table of attribute-values and probabilities reflecting the adopted instances (see Figure 1).
Figure 1: A new instance is incorporated into an empty Cobweb tree. Cobweb adds it to the root and updates the root’s count table to reflect the instance’s attribute-values.
@@
During learning, Cobweb categorizes new instances down its tree recursively and updates the count tables of concepts along the categorization path to reflect the attribute values of the instances (Figure 2(a)). At each branch point, Cobweb considers four restructuring operations: adding (adding the instances to the best child concept and updating the attribute value counts), merging (combining the two “best” children and reconsidering the available operations), splitting (removing the “best” child and elevating its children to the current level), and creating (creating a new child node which adopts the new learned instance only) (see Figure 2(b) from MacLellan et al. (2016) ).
During prediction, Cobweb follows a similar approach to learning, but it does not modify the counts of concepts along the path. Given an instance (potentially missing some of its attribute values), Cobweb sorts it downward through its hierarchy: Starting at the root node, at every branch (a concept node with its children), Cobweb recursively considers whether to direct the instance to one of the children or whether to stop at the current node. When this process terminates, it uses the final concept’s count table to predict the values of unobserved attributes.
During both prediction and learning, Cobweb uses a measure called category utility (Corter & Gluck, 1992) to evaluate its choices, and proceeds with the operation that maximizes category utility. Category utility evaluates the increase in the predictive power of a child node compared to its parent, and is similar to the information gain metric used in decision tree induction (Quinlan, 1986); however, instead of supporting prediction of a single attribute, it has the capacity to support prediction of all attributes. The function is defined as:
$\frac{\sum_{k=1}^{n}P(C_{k})\left[\sum_{i}\sum_{j}P(A_{i}=V_{ij}|C_{k})^{2}-P(A_{i}=V_{ij})^{2}\right]}{n}$ (1)
Where $n$ is the number of concepts, $P(C_{k})$ is the overall probability of the $k$th child concept, $P(A_{i}=V_{ij}|C_{k})$ is the probability of attribute $A_{i}$ having value $V_{ij}$ given concept $C_{k}$, and $P(A_{i}=V_{ij})$ is the probability of attribute $A_{i}$ having value $V_{ij}$ in the parent concept. In general terms, $\sum_{i}\sum_{j}P(A_{i}=V_{ij}|C_{k})^{2}$ represents the expected number of attributes correctly predicted for a given child $C_{k}$, while $\sum_{i}\sum_{j}P(A_{i}=V_{ij})^{2}$ corresponds to the expected number of correct guesses made from the parent. This equation represents the average increase in the number of correctly guessed attribute-values within the children compared to the parent, weighted by the probability of each child. To compare cases with different numbers of children, this score is normalized by dividing by $n$, the number of children.
While the original Cobweb approach only supports nominal attributes, Cobweb/3 (McKusick & Thompson, 1990) extends this support to continuous attributes by using normal distributions to describe their probability densities. Each concept stores the mean and standard deviation for each continuous attribute rather than the counts that are maintained for nominal attributes (see Figure 3).
(a) Cobweb categorizes a new instance down its hierarchy and updates the attribute value counts.
(b) Four update operations that Cobweb considers at each node during learning (MacLellan et al., 2016).
Figure 2: Cobweb’s Learning Process
3 Cobweb/4V
@@
This paper introduces Cobweb/4V, a new Cobweb variant designed for visual learning tasks. It extends a prior implementation by MacLellan et al. (2016) with a new prediction approach that uses multiple concepts throughout the categorization tree to yield improved predictive performance. It also introduces a new tensor-based concept representation that enables efficient processing of images.
3.1 Learning with Mutual Information
@@
All prior Cobweb studies employed the probability-theoretic category utility (Corter & Gluck, 1992) in both predicting and learning processes, which calculates the expected increase in the ability to predict attribute values correctly given the knowledge of an instance’s membership in a particular concept. An analysis by Fisher (1996) suggests this formulation is an unsupervised extension of the Gini Index measure commonly used to construct decision trees. In Cobweb/4V, however, we use the information-theoretic category utility (Corter & Gluck, 1992). This involves establishing a connection between feature predictability and informativeness. Using this new measure yields the following updated category utility function:
$\frac{\sum_{k=1}^{n}P(C_{k})\left[H(A=V)-H(A=V|C_{k})\right]}{n}$ (2)
where $H(A=V)=\sum_{i}\sum_{j}[-P(A_{i}=V_{ij})\log(P(A_{i}=V_{ij})]$ is the entropy over all attribute values in the parent and $H(A=V|C_{k})$ is the entropy over all attribute values in the given child $k$. This measure provides an unsupervised extension of the information gain measure commonly used in decision trees. It is analogous to mutual information, evaluating the increase in the information known about the attribute values gained by knowledge of the children over the parent. As the Gini index is a first-order approximation of information gain, this new variation of category utility supports higher precision than the probability-theoretic variation. Also, by formulating category utility in terms of entropy, we can easily use different distributions to represent attributes, as many commonly used distributions have closed-form formulas for computing their entropy.
3.2 Predicting with a Combination of Concepts
@@
During prediction, Cobweb first finds a final concept at the subordinate level in a greedy way, then uses the count information associated with this node to predict the unobserved attribute values of the categorized instance. This approach has been widely adopted in prior studies because it seems to yield good empirical performance (MacLellan et al., 2016; MacLellan & Thakur, 2022; MacLellan et al., 2022). However, other studies have explored alternative concept levels for predictions, such as using basic-level concepts (Fisher & Langley, 1990; Corter & Gluck, 1992).
@@
Our revised approach introduces a new prediction strategy that combines predictions from multiple concepts within the tree. Given an instance $x$ with some unobserved feature(s) and a maximum number of the concepts to use for prediction ($N_{max}$), Cobweb begins categorization at the root. However, instead of sorting $x$ down a single, greedy path, Cobweb/4V instead uses best-first search. At each point in the search, it expands the node $c^{*}$ on the search frontier that has the greatest score $s(c)=P(c|x)P(x|c)$. This heuristic, which is sometimes referred to as collocation (Jones, 1983), [^1] Prior work shows it is related to probabilistic category utility; the expected correct guesses of a concept can be reinterpreted as a form of weighted collocation (Corter & Gluck, 1992). is the product of cue and category validity. Once a $c^{*}$ is identified, the system adds it to an expanded node list. Next, the system evaluates the collocation of the children of $c^{*}$ and adds these to the search frontier to be considered for expansion. This process repeats until Cobweb/4V has expanded $N_{max}$ nodes. If $\mathcal{C}^{*}$ denotes all the nodes expanded during categorization, Cobweb/4V predicts the attribute $X_{i}$ will have value $x_{i}$ with probability:
@@
This represents the combination of predictions from all expanded nodes, weighted by their collocation (the softmax ensures the sum of the weights to one). Although Cobweb/4V only expands $N_{max}$ nodes, rather than all the nodes, this approach is effectively a form of Bayesian model averaging (Hinne et al., 2020).
Figure 3: The single image on the left shows the tensor-based representation for instances. The concepts store the means and standard deviations for the pixel attributes in internal tensor representations as well.
3.3 A New Tensor Representation
In Cobweb/4V, each instance is represented by a tensor of pixels and a class label, as opposed to a list of attribute-value pairs, as used in previous Cobweb approaches. Given our focus on vision tasks, this adaptation better enables Cobweb/4V to support vision learning tasks, where each input takes the form of an $n$-channel 2D image, which is represented as a matrix of pixel intensity values. [^2] We use grayscale images in our experiments, so the inputs are just a single channel (i.e., $n=1$), but the approach supports multi-channel input as well. Each instance is also paired with a nominal digit label, which Cobweb/4V natively supports as a categorical, rather than requiring a specialized encoding, such as one-hot. Similar to Cobweb/3, each node stores the means and standard deviations of the pixel features for all the images that have been categorized under it. However, it uses a tensor representation to store this statistics, which enables more efficient inference and updating. Also, like Cobweb, it stores digit labels in a probability table that keeps track of how frequently each label appears across instances assigned to a given concept (Figure 3).The values for each attribute are assumed to be conditionally independent given the concept. Consequently, the uncertainty of a node $c$ is the sum of entropy values for the attributes under $c$. The use of tensor representations for instances and concepts (our implementation uses PyTorch) significantly enhances speed, yielding substantially faster processing than prior versions of Cobweb that use attribute-value lists.
4 Experiments
We evaluate Cobweb/4V and two neural network baselines across distinct experimental conditions with the MNIST dataset (60,000 images in the training set, 10,000 images in the test set), prevalently used in the continual learning studies (van de Ven et al., 2022; LeCun, 1998). [^3] The codes for the experiments are available at https://github.com/Teachable-AI-Lab/cobweb-vision
4.1 Preliminary Experiment